OpenStudy (anonymous):

A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical? i got 0.5?

OpenStudy (saifoo.khan):

u should get 2.5

OpenStudy (anonymous):

i though the answer was 5

OpenStudy (saifoo.khan):

nope!

OpenStudy (anonymous):

well heres what i did x=amount of 10% solution used 10x+40(10-x)=25(10) x+4(10-x)=25(1) x+40-4x=25 3x=15 so x=5

OpenStudy (saifoo.khan):

.10X+.40(10-X)=.25*10 .10X+4-.4X=2.5 .3X=2.5-4 .3X=1.5 X=1.5/.3 X=5 OZ. OF EACH MIXTURE IS NEEDED. PROOF: .10*5+.40*5=2.5 .5+2=2.5 2.5=2.5

OpenStudy (anonymous):

the question asks how much of each mixture and that is what x stands for so the answer would be 5, no?